OpenStudy (anonymous):

Check my answers please? Only 5 multiple choice questions. I'm really confused. Any and all help is appreciated! Will give medal! 1. Which of the following is a correctly written thermochemical equation? NH4Cl NH4+ + Cl 2C8H18 + 25O2 16CO2 + 18H2O, H = 5,471 kJ/mol Fe (s) + O2 (g) Fe2O3 (s), H = 3,926 kJ -->C3H8 (g) + 5O2 (g) 3CO2 (g) + 4H2O (l), H = 2,220 kJ/mol 2. How does a phase change affect a thermochemical equation? It alters the moles of reactants. -->It can affect the (delta) H value. It alters the products. It affects the balance of the equation.

OpenStudy (anonymous):

3. For the reaction CH4 (g) + 2O2 (g) CO2 (g) + 2H2O (l), H = 890 kJ, what will be the change in enthalpy when 3 moles of methane react in excess oxygen? 890 kJ -->2.67 × 103 kJ +890 kJ +2.67 × 103 kJ 4. In which of the following thermochemical equations would the H be considered a heat of solution? NH4Cl (s) NH4+ (aq) + Cl (aq), H = + 14.8 kJ/mol 2C8H18 (l) + 25O2 (g) 16CO2 (g) + 18H2O (l), H = 5,471 kJ/mol -->4Fe (s) + 3O2 (g) 2Fe2O3 (s), H = 3,926 kJ C3H8 (g) + 5O2 (g) 3CO2 (g) + 4H2O (l), H = 2,220 kJ/mol 5. For the following reaction at STP, how much heat will be required to produce 63 L of NO? (At STP, one mole of any gas occupies 22.4 L.) N2 (g) + O2 (g) 2NO (g), H = 180.5 kJ +63 kJ +127 kJ -->+254 kJ +508 kJ

OpenStudy (anonymous):

Are they alright?

OpenStudy (anonymous):

All but number 4 are correct, it should be NH4Cl (s) NH4+ (aq) + Cl (aq), H = + 14.8 kJ/mol

OpenStudy (anonymous):

Alright, thanks for responding...I'm long done with this by now but thank you anyways! :) I appreciate it!